OpenStudy (anonymous):

a block of wood was cut into 3. the 3 blocks have a ratio of 3:4:5 if the middle brick has a length of 10, what was the size of the original block of wood?

OpenStudy (anonymous):

please i need this help me

OpenStudy (anonymous):

since the ratio is 3:4:5 then let their respective lengths be 3x,4x,5x given the middle brick is of length 10 thus by theprob we have 4x=10 or x=10/4=2.5 hence the length of first brick is 3x =3*.5 =7.5 and that of the third brick =5x =5*2.5 =12.5